# From Concept to Reality: 5G Positioning with Open-Source Implementation of UL-TDoA in OpenAirInterface

### v2

@@ I Introduction
The introduction of 5G New Radio (NR) technology represents a significant evolution in wireless communications, bringing unparalleled advancements in speed, connectivity, and latency. Beyond merely enhancing existing mobile broadband services, 5G NR is expected to transform numerous industries by enabling ultra-reliable low-latency communication (URLLC), massive machine-type communications (mMTC), and enhanced mobile broadband (eMBB). A key innovation within the 5G framework is precise positioning and localization, which is anticipated to create new opportunities across sectors such as automotive, healthcare, logistics, and smart city development.
@@ III-A UL-TDoA Positioning Procedure in OAI
The UL-TDoA positioning procedure begins with an external API generating a location information request for a target UE for the LMF. As shown in Figure  5 to initiate the location procedure, the external API (which can be an advanced API or a simple one-line command) sends the HTTP post request to determine-location API of LMF, where the request contains a data structure of type InputData [10, Section 6.1.6.2.2]. To enable the exchange of this message, we first developed the 3gpp-compliant LMF framework and integrated it into OAI’s core network framework. Then within that LMF framework, we develop the determine-location API [10, Section 6] to handle the positioning request following the 3gpp standard.
@@ III-B Enabling ToA estimation in OAI’s gNB
$\hat{\mathbf{H}}_{n,p,l}^{\text{interp}}(k)\!\!=\!\!\begin{cases}\!\mathbf{F}_{\text{start}}\cdot\hat{\mathbf{H}}_{n,p,l}[k]&\!\!\!\!\text{if }\!k=0\text{ or }k<K_{\text{TC}}\\ \!\mathbf{F}_{\text{middle}}\cdot\hat{\mathbf{H}}_{n,p,l}[k]&\!\!\!\!\text{if }\!k\neq 0\text{ and }(k+K_{\text{TC}})<N_{\text{sc}}\\ \!\mathbf{F}_{\text{end}}\cdot\hat{\mathbf{H}}_{n,p,l}[k]&\!\!\!\!\text{if }\!(k+K_{\text{TC}})\!\geq\!N_{\text{sc}}\!\text{ or }\!k\!=\!M_{\text{sc}}\!-\!1\end{cases}$ (2)
@@
$\hat{\mathbf{H}}_{n,p,l}^{\text{oversamp}}[k]=\begin{bmatrix}\hat{\mathbf{H}}_{n,p,l}^{\text{interp}}(1:N_{\text{fft}}/2),\\ \mathbf{zeros}(1,L\times N_{\text{fft}}-N_{\text{fft}}),\\ \hat{\mathbf{H}}_{n,p,l}^{\text{interp}}(N_{\text{fft}}/2+1:\text{end})\end{bmatrix}$ (3)
@@
$\hat{\mathbf{h}}^{\text{oversamp}}_{n,p,l}(t)=\frac{1}{L\times N_{\text{fft}}}\sum_{k=0}^{L\times(N_{\text{fft}}-1)}\hat{\mathbf{H}}_{n,p,l}^{\text{oversamp}}[k]e^{i2\pi kt/L\times N_{\text{fft}}}$ (4)
@@
The FAPI interface is a standardized interface between the PHY and the MAC specified by the small cell forum [13]. The SRS measurements are sent from PHY to the MAC via the SRS.indication message. This message includes a field called “Timing advance offset in nanoseconds,” which is meant to be used for UL TDoA positioning measurements. However, the standard does not foresee that there can be multiple such measurements from multiple TRPs connected to the same DU. We solved this issue by adding a new value for SRS type (5) and a new report type (“Localization”) to the SRS indication message. These additions are allowed in the current specification and do not break the compatibility with implementations that do not support this report type. The actual values of the Timing advance offsets in ns from the multiple TRPs are sent in the Report tagged list value (TLV) directly as an array in the value field, where each value uses 16bit.
@@ IV-B3 Ground truth measurements
IV-C Results
This section describes the process of validating the overall performance of a deployed 5G network by testing its end-to-end functionality. The validation involves running the OAI 5G CN components, including the oai-lmf, alongside a gNB that integrates CU-DU functions, and a Firecell’s ORAN RU. A 5G phone, acting as a UE, is then registered with this network, and the success of this registration is verified by analyzing the logs generated by the oai-amf as shown in Figure 10. Once we ensure the UE and gNB are connected we trigger the determine location API to start the positioning request.
We anticipate receiving the relative Cartesian location from the gNB configuration file and the uL-RTOA values from the corresponding reported TRPs at the LMF. These inputs will be utilized in the position estimation function. The LMF is running a Least Squares estimation algorithm for a coarse estimation that will be then used in a Non Linear Least Squares estimation method akin to [17] for a more fined grained positioning. The evaluation of this process is reflected in the oai-lmf logs as follows:
[lmf_app] [info] handle measurement response
gnbId:3,trpId:1,trpRelCartLoc(x:0,y:0,z:2),k1:563
gnbId:3,trpId:2,trpRelCartLoc(x:-9,y:0,z:2),k1:563
gnbId:3,trpId:3,trpRelCartLoc(x:-27,y:0,z:2),k1:563
gnbId:3,trpId:4,trpRelCartLoc(x:-36,y:0,z:2),k1:563
ToA Values:5.72717e-07
ToA Values:5.72717e-07
ToA Values:5.72717e-07
ToA Values:5.72717e-07
Estimated Position: x = 25.4048, y = 11.8080
And finally the estimated position is returned to the determine-location API.
Point Error (m) Point Error (m)
A 4.7766 I 3.0248
B 4.6982 J 2.3258
C 0.8191 K 3.3805
D 1.3082 L 3.2353
E 2.3743 M 0.3461
F 2.0003 N 4.8525
G 2.2792 O 5.4106
H 3.3180 P 4.1295
TABLE II: MAE Error associated with Points A-P
@@ V Conclusions
This paper provided the first open-source implementation of the 3GPP Uplink Time Difference of Arrival (UL-TDoA) positioning method within the OpenAirInterface (OAI) framework, successfully integrating UL-TDoA into both the RAN and CN components. By adhering to 3GPP standards, this implementation enabled precise and real-time positioning of user equipment (UE) in 5G networks, offering a flexible alternative to proprietary solutions. The approach was validated through both simulation and real-world testing, demonstrating its reliability and accuracy. This work not only enhanced the capabilities of OAI for 5G positioning but also contributed to the broader research community by providing a valuable tool for further innovation and collaboration in the field of cellular network positioning technologies.

### v3

@@ I Introduction
5G new radio (NR) represents a major leap in wireless communications, significantly improving speed, connectivity, and latency. Beyond enhancing mobile broadband (eMBB), 5G NR enables ultra-reliable low-latency communication (URLLC) and massive machine-type communications (mMTC). A key innovation within the 5G framework is precise positioning and localization, which is anticipated to create new opportunities across sectors such as automotive, healthcare, logistics, and smart city development.
@@ III-A UL-TDoA Positioning Procedure in OAI
The UL-TDoA positioning procedure begins with an external API generating a location information request for a target UE for the LMF. As shown in Figure  5 to initiate the location procedure, the external API (which can be an advanced API or a simple one-line command) sends the HTTP post request to determine-location API of LMF, where the request contains a data structure of type InputData [10, Section 6.1.6.2.2]. To enable the exchange of this message, we first developed the 3gpp-compliant LMF framework and integrated it into OAI’s core network framework[12]. Then within that LMF framework, we develop the determine-location API [10, Section 6] to handle the positioning request following the 3gpp standard.
@@ III-B Enabling ToA estimation in OAI’s gNB
$\hat{\mathbf{H}}_{n,p,l}^{\text{interp}}[k]\!\!=\!\!\begin{cases}\!\mathbf{F}_{\text{start}}\cdot\hat{\mathbf{H}}_{n,p,l}[k]&\!\!\!\!\text{if }\!k=0\text{ or }k<K_{\text{TC}}\\ \!\mathbf{F}_{\text{middle}}\cdot\hat{\mathbf{H}}_{n,p,l}[k]&\!\!\!\!\text{if }\!k\neq 0\text{ and }(k+K_{\text{TC}})<N_{\text{sc}}\\ \!\mathbf{F}_{\text{end}}\cdot\hat{\mathbf{H}}_{n,p,l}[k]&\!\!\!\!\text{if }\!(k+K_{\text{TC}})\!\geq\!N_{\text{sc}}\!\text{ or }\!k\!=\!M_{\text{sc}}\!-\!1\end{cases}$ (2)
@@
$\hat{\mathbf{H}}_{n,p,l}^{\text{oversamp}}=\begin{bmatrix}\hat{\mathbf{H}}_{n,p,l}^{\text{interp}}(1:N_{\text{fft}}/2),\\ \mathbf{zeros}(1,L\times N_{\text{fft}}-N_{\text{fft}}),\\ \hat{\mathbf{H}}_{n,p,l}^{\text{interp}}(N_{\text{fft}}/2+1:\text{end})\end{bmatrix}$ (3)
@@
$\hat{\mathbf{h}}^{\text{oversamp}}_{n,p,l}(t)=\frac{1}{L\times N_{\text{fft}}}\sum_{k=0}^{L\times N_{\text{fft}}-1}\hat{\mathbf{H}}_{n,p,l}^{\text{oversamp}}[k]e^{i2\pi kt/(L\times N_{\text{fft}})}$ (4)
@@
The FAPI interface is a standardized interface between the PHY and the MAC specified by the small cell forum [14]. The SRS measurements are sent from PHY to the MAC via the SRS.indication message. This message includes a field called “Timing advance offset in nanoseconds,” which is meant to be used for UL TDoA positioning measurements. However, the standard does not foresee that there can be multiple such measurements from multiple TRPs connected to the same DU. We solved this issue by adding a new value for SRS type (5) and a new report type (“Localization”) to the SRS indication message. These additions are allowed in the current specification and do not break the compatibility with implementations that do not support this report type. The actual values of the Timing advance offsets in ns from the multiple TRPs are sent in the report-tagged list value (TLV) directly as an array in the value field, where each value uses 16bit.
@@ IV-B3 Ground truth measurements
IV-C Preprocessing and Position Estimation
In a multi-RU positioning setup, precise RU synchronization is crucial for accurate TDoA measurements, as just a 1-nanosecond error can cause a 0.3-meter positioning inaccuracy. While in some setups this accuracy can be achieved, in our GEO5G testbed we have observed timing errors of up to 40ns. Using a common reference TRP across all RUs will therefore result in inaccurate TDoA measurements and degraded positioning estimates.
To mitigate this, a per-RU reference approach is adopted, where within each RU, one TRP is designated as the local reference for TDoA calculations, ensuring that all measurements within the RU remain internally consistent and unaffected by synchronization drifts in other RUs.
Next, to ensure robust TDoA estimation while mitigating synchronization errors, a two-stage filtering mechanism is employed. Firstly, by averaging over the $N$ most recent TDoAs from each RU, we avoid small-scale fluctuations in the measurements. Secondly, for compensating larger spikes in the TDoAs, another filter is used that exploits the geographical information of the testing area. Knowing the minimum and maximum possible TDoA values based on known TRP positions, we define a filter that discards any invalid TDoA measurement and retains only those within a predefined bound, along with their corresponding TRP coordinates. This filtering approach effectively mitigates the impact of noisy measurements caused by multipath, Non-Line-of-Sight (NLoS) conditions, and hardware impairments.
IV-D Results
This section validates the performance of our deployed 5G network by registering a UE and triggering the location determination API. The TRP coordinates and UL-RToA values from each TRP are used in a stochastic optimization method for position estimation from our previous work [18]. The positioning error was evaluated for two setups: a single RU (RU3) and a multi-RU (RU2 and RU3). The results show that multiple RUs improve accuracy by increasing TRP diversity in both the x and y axes. In the single RU setup, the TRPs are arranged linearly, causing higher uncertainty in the y-axis, which is reduced with the multi-RU setup. Points A, B, O, and P on the edge of the testing area show degraded accuracy due to NLoS, multipath and diffraction conditions to the buildings around them. Table II, is summarizing the Mean Absolute Error (MAE) for both setups. Also, Figure 18 shows a mobile scenario where a person holds the UE and takes a trajectory covering all testing points in the multi-RU setup.
Point 1 RU MAE (m) 2 RUs MAE (m) Point 1 RU MAE (m) 2 RUs MAE (m)
A 4.77 1.88 I 3.02 0.60
B 4.69 1.13 J 2.32 0.83
C 0.81 0.82 K 3.38 0.82
D 1.30 0.83 L 3.23 1.20
E 2.37 0.95 M 3.34 0.64
F 2.00 0.55 N 4.85 0.68
G 2.27 0.92 O 5.41 0.97
H 3.31 0.56 P 4.12 1.98
TABLE II: MAE for single RU (RU3) and multi-RU (RU2+RU3) setups at points A-P
In future developments, a multi-gNB setup with additional RUs and increased TRPs is expected to enhance accuracy by enabling techniques such as multipath and NLoS mitigation.
Figure 18: UE tracking in a mobile scenario
@@ V Conclusions
This paper provided the first open-source implementation of the 3GPP UL-TDoA positioning method within the OAI framework, successfully integrating UL-TDoA into both the RAN and CN components. By adhering to 3GPP standards, this implementation enabled precise and real-time positioning of UE in 5G networks, offering a flexible alternative to proprietary solutions. The approach was validated through both simulation and real-world testing, demonstrating its reliability and accuracy. This work not only enhanced the capabilities of OAI for 5G positioning but also contributed to the broader research community by providing a valuable tool for further innovation and collaboration in the field of cellular network positioning technologies.